# Reduced Order Modelling, Dynamical Systems, Deep Learning, Uncertainty Quantification journal: Journal name

### v1

@@ Abstract
UP-ROM : Uncertainty-Aware and Parametrised dynamic Reduced-Order Model – application to unsteady flows
@@ 1 Introduction
Through UP-ROM, we aim to contribute to these advances by proposing a Transformer-based, dynamic, parametric and variational ROM. This model is designed to solve unsteady PDEs while considering and adapting to external excitation variables through cross-attention mechanisms, thus exploiting breakthroughs in large language models (LLMs) for PDE applications. In addition, it provides a probabilistic counterpart to most deterministic dynamic modelling baselines, enabling uncertainty quantification. Crucially, our approach achieves this in a computationally efficient manner, significantly reducing the cost compared to Gaussian processes or Bayesian models, while maintaining robustness and scalability for complex dynamical systems.
@@ 2.2 Temporal evolution
Figure 1: UP-ROM architecture with $q$ the look-back window
@@ 2.3 Model training
where $\mathcal{E}$ is the encoding process, $\mathcal{D}$ the decoding process, KLD the Kullback-Leibler divergence and $\hat{z}$ the latent variables. In addition, $\lambda$ and $\beta$ are regularisation terms. The first term of this loss promotes a mirroring process between encoder and decoder, similar to a classical projection. The Kullback-Leibner divergence term is related to the use of the VAE and is explained in section 4.
2.4 Validation of UP-ROM – application to flow around a cylinder
@@
Figure 2: Predicted kinetic energy signal using UP-ROM compared with the ground truth.
@@
Figure 4: UP-ROM parametrised by external excitation variable $\xi$.
@@
To test whether the performance of UP-ROM extends to a larger dimensional parameter space, the aspect ratio of the bluff body was added as a second parameter. The aspect-ratio, as defined in Figure 6, characterizes the shape of the bluff body and significantly influences the flow behaviour. An aspect-ratio smaller than 1 indicates that the obstacle is taller than wide, creating a stronger disruption in the flow. This increased blockage leads to earlier vortex shedding and a bifurcation occurring at a lower Reynolds number. In contrast, an aspect-ratio greater than 1 corresponds to an elongated body, similar to an airfoil, which disturbs the flow less significantly. In this case, the onset of vortex shedding and bifurcation is delayed, occurring at a higher Reynolds number due to the more streamlined geometry. This effect is illustrated in Figure 6, where different aspect ratios result in distinct flow regimes, even though the Reynolds number remains unchanged.
@@
Performance with a two-variable parametrisation $(R_{e},\beta)$ can be visualised using a pointwise scaled MSE, as shown in Figure 7, which ensures that each grid point is equally weighted in the evaluation. The scaled MSE, computed on test data, is obtained locally using the magnitude of local variations. The resulting heatmap provides a a-posteriori perspective on the model’s ability to generalise and its overall performance. The signals shown on the map represent the evolution of the kinetic energy for that configuration over time. The model performs well on the training sets, demonstrating its ability to learn and predict. In particular, its interpolation results are strong. However, similar behaviour to the single parameter case is observed. UP-ROM is able to distinguish between stable and limit-cycle dynamics in the two-parameter regime. However, the accuracy of the prediction deteriorates when the model is used in extrapolation mode.
@@ 4 Uncertainty quantification
Most conventional reduction strategies rely on deterministic estimation, where the latent space projection is represented by a single point estimate. This approach fails to capture the inherent uncertainty within the model. In contrast, UP-ROM utilizes variational methods to assess uncertainties, providing a probabilistic projection in the latent space instead of a deterministic one. This is achieved through the use of a Variational Autoencoder (VAE). Specifically, given the observed data $X=D$ (i.e., the data set $D$), the method seeks to infer the posterior distribution $p(z|X=D)$, where $z$ represents the latent variables.
@@ A Variational Autoencoder (VAE) integrates the Evidence Lower Bound (ELBO) into its loss function, as
Figure 8: UP-ROM Uncertainty Quantification process after inference
@@ 5 Adaptive Sampling for Optimized Generalization
After completing five retraining iterations, the UP-ROM method demonstrates consistent and uniform learning performance across the entire range of model parameters, effectively capturing the underlying bifurcation dynamics. Notably, the model’s uncertainty appears to have converged to a stable final distribution, denoted as $\nu_{\xi}$. The model progressively identifies parameter points that minimize the error at each training step. As retraining continues at these optimal points, we observe a convergence of both the scaled mean squared error (MSE) and the uncertainty, which both tend toward an optimal distribution.
@@ 6 Conclusion
This paper presents a robust and parametrised reduced-order model (ROM) for time-dependent partial differential equations applied to the Navier-Stokes equations in the laminar regime. The model makes use of memory embedding and variational inference, first using a nonlinear variational autoencoder to reduce the high-dimensional dynamics to a low-dimensional latent space, and second using an attention-based transformer to capture time dependencies in the reduced manifold. In addition, cross-attention incorporates dependencies on external parameters. This allows UP-ROM to handle different dynamical regimes, making it adaptable to external excitation variables and enhancing generalisation.

### v2

@@ Abstract
UP-dROM : Uncertainty-Aware and Parametrised dynamic Reduced-Order Model – application to unsteady flows
@@ 1 Introduction
Through UP-dROM, we aim to contribute to these advances by proposing a Transformer-based, dynamic, parametric and variational ROM. This model is designed to solve unsteady PDEs while considering and adapting to external excitation variables through cross-attention mechanisms, thus exploiting breakthroughs in large language models (LLMs) for PDE applications. In addition, it provides a probabilistic counterpart to most deterministic dynamic modelling baselines, enabling uncertainty quantification. Crucially, our approach achieves this in a computationally efficient manner, significantly reducing the cost compared to Gaussian processes or Bayesian models, while maintaining robustness and scalability for complex dynamical systems.
@@ 2.2 Temporal evolution
Figure 1: UP-dROM architecture with $q$ the look-back window
@@ 2.3 Model training
where $\mathcal{E}$ is the encoding process, $\mathcal{D}$ the decoding process, KLD the Kullback-Leibler divergence and $\hat{z}$ the latent variables. In addition, $\lambda$ and $\beta$ are regularisation terms. The first term of this loss promotes a mirroring process between encoder and decoder, similar to a classical projection. The Kullback-Leibler divergence term is related to the use of the VAE and is explained in section 4.
2.4 Validation of UP-dROM – application to flow around a cylinder
@@
Figure 2: Predicted kinetic energy signal using UP-dROM compared with the ground truth.
@@
Figure 4: UP-dROM parametrised by external excitation variable $\xi$.
@@
To test whether the performance of UP-dROM extends to a larger dimensional parameter space, the aspect ratio of the bluff body was added as a second parameter. The aspect-ratio, as defined in Figure 6, characterizes the shape of the bluff body and significantly influences the flow behaviour. An aspect-ratio smaller than 1 indicates that the obstacle is taller than wide, creating a stronger disruption in the flow. This increased blockage leads to earlier vortex shedding and a bifurcation occurring at a lower Reynolds number. In contrast, an aspect-ratio greater than 1 corresponds to an elongated body, similar to an airfoil, which disturbs the flow less significantly. In this case, the onset of vortex shedding and bifurcation is delayed, occurring at a higher Reynolds number due to the more streamlined geometry. This effect is illustrated in Figure 6, where different aspect ratios result in distinct flow regimes, even though the Reynolds number remains unchanged.
@@
Performance with a two-variable parametrisation $(R_{e},\beta)$ can be visualised using a pointwise scaled MSE, as shown in Figure 7, which ensures that each grid point is equally weighted in the evaluation. The scaled MSE, computed on test data, is obtained locally using the magnitude of local variations. The resulting heatmap provides a a-posteriori perspective on the model’s ability to generalise and its overall performance. The signals shown on the map represent the evolution of the kinetic energy for that configuration over time. The model performs well on the training sets, demonstrating its ability to learn and predict. In particular, its interpolation results are strong. However, similar behaviour to the single parameter case is observed. UP-dROM is able to distinguish between stable and limit-cycle dynamics in the two-parameter regime. However, the accuracy of the prediction deteriorates when the model is used in extrapolation mode.
@@ 4 Uncertainty quantification
Most conventional reduction strategies rely on deterministic estimation, where the latent space projection is represented by a single point estimate. This approach fails to capture the inherent uncertainty within the model. In contrast, UP-dROM utilizes variational methods to assess uncertainties, providing a probabilistic projection in the latent space instead of a deterministic one. This is achieved through the use of a Variational Autoencoder (VAE). Specifically, given the observed data $X=D$ (i.e., the data set $D$), the method seeks to infer the posterior distribution $p(z|X=D)$, where $z$ represents the latent variables.
@@ A Variational Autoencoder (VAE) integrates the Evidence Lower Bound (ELBO) into its loss function, as
Figure 8: UP-dROM Uncertainty Quantification process after inference
@@ 5 Adaptive Sampling for Optimized Generalization
After completing five retraining iterations, the UP-dROM method demonstrates consistent and uniform learning performance across the entire range of model parameters, effectively capturing the underlying bifurcation dynamics. Notably, the model’s uncertainty appears to have converged to a stable final distribution, denoted as $\nu_{\xi}$. The model progressively identifies parameter points that minimize the error at each training step. As retraining continues at these optimal points, we observe a convergence of both the scaled mean squared error (MSE) and the uncertainty, which both tend toward an optimal distribution.
@@ 6 Conclusion
This paper presents a robust and parametrised reduced-order model (ROM) for time-dependent partial differential equations applied to the Navier-Stokes equations in the laminar regime. The model makes use of memory embedding and variational inference, first using a nonlinear variational autoencoder to reduce the high-dimensional dynamics to a low-dimensional latent space, and second using an attention-based transformer to capture time dependencies in the reduced manifold. In addition, cross-attention incorporates dependencies on external parameters. This allows UP-dROM to handle different dynamical regimes, making it adaptable to external excitation variables and enhancing generalisation.